# Predictive Coding Networks and Inference Learning: Tutorial and Survey

### v1

@@ 3.2.2. Multi-layer PC
As in the simple model above, we assume a multidimensional Gaussian for the conditional distribution of each layer, as well as for the prior: [^10] This is a reasonable assumption for traditional (fully-connected, feedforward) ANNs with large layer widths, $n_{\ell}\sim n\gg 1$, which becomes exact in the so-called large-width limit $n\mapsto\infty$, due to the central limit theorem. However, non-Gaussianities appear at finite width (see, e.g., (roberts_principles_2022, ; grosvenor_edge_2022, ) for theoretical treatments), and it is unclear how realistic this assumption is in more general networks involving intralayer or feedback connections.
@@
which is the energy of Section 2; cf. also (27). Applying EM to this then yields IL as presented in section 2.
@@ 3.3.1. Discriminative and Generative PCNs
At this point we can provide a number of interesting interpretations of earlier sections. Notice that in the derivation above, the model had observed variables at layer $\ell=0$. This corresponds to generative PCN, as discussed in section 2.5. If trained in supervised mode, we can understand the clamping of the label $\bm{y}^{(n)}$ to the final layer $\bm{z}^{p}$ as follows: $\bm{y}^{(n)}$ becomes a parameter of the prior (e.g. its mean). If trained in unsupervised mode, it is treated just like the other hidden variables. Interestingly then, for discriminative PCN, we see that the clamped data $\bm{x}^{(n)}$ is a parameter for the prior, and the label is now the observed variable. This is shown in table 6.
Table 6. Generative and discriminative models together with learning modes in the probabilistic perspective of PCNs.
@@ 3.3.2. Precision Matrices
In the context of predictive coding, the covariance matrices $\Sigma^{\ell}$ are often interpreted as the uncertainty at each layer around the mean $\bm{\mu}^{\ell}$, with their inverse $(\Sigma^{\ell})^{-1}\eqqcolon\Pi^{\ell}$ sometimes called precision matrices (friston_learning_2003, ; millidge_theoretical_2022, ; marino_predictive_2022, ). Within neuroscience and psychology, they have been used in a wide range of models (millidge_predictive_2022, ), where it has been suggested that they implement a type of attention (feldman_attention_2010, ). By up-weighting the error, increasing the precision of a variable would serve as a form of gain modulation. From the ML perspective, instead of setting them to identity matrices, one can add minimization of $E$ w.r.t. $\Pi^{\ell}$ as an additional part of the M-step, seeing them simply as additional parameters of the generative model. However, their practical utility remains unclear, with varying results. In particular, (ofner_generalized_2022, ) do not find advantages of additionally learning for generative modeling tasks, while (ororbia_neural_2022, ) do appear to find some benefit, suggesting they act as a type of lateral modulation that induces a useful form of sparsity in the model. Comparisons with VAEs are made in (pinchetti_predictive_2022, ), with a different theoretical interpretation by (millidge_theoretical_2022, ) (see section 4 for further discussion).
@@ 3.4. Connections to Other Latent Variable Models
With $f$ as a linear function, the two-layer PCN model discussed above is equivalent to a linear factor model. This is illustrated in fig. 13. Specifically, if we take $\bm{\mu}_{z}=\bm{0}$, $\Sigma_{z}=I$, and $\Sigma_{x}=\operatorname{diag}\bm{\sigma}^{2}$ with $\bm{\sigma}^{2}=[\sigma_{1}^{2},\sigma_{2}^{2},...,\sigma_{n}^{2}]^{T}$, then we obtain exactly factor analysis. If additionally take $\Sigma_{x}=\sigma^{2}I$, i.e. the variances $\sigma_{i}^{2}$ are all equal to each other, we obtain probabilistic PCA (goodfellow_deep_2016, ). Like in PCNs, parameters in this model can be learned using the EM algorithm (bishop_pattern_2006, ).
Figure 13. Graphical model comparison of PCNs and other probabilistic models, with $\theta$ model parameters and $\phi$ variational parameters (cf. appendix A). Standard PCNs (left column) use MAP inference of the latent variables during the E-step, without variational parameters. Using a Gaussian variational (middle column) adds the mean and variance of the variational as parameters $\phi$ which may additionally be varied, shown by dotted lines. This is similar to $\phi$ in a VAE (top right). VAEs differ from PCNs in that their variational posterior is conditioned explicitly on $\bm{x}$ (i.e. $q(\bm{z}|\bm{x})$ in VAEs vs. $q(\bm{z})$ in PCNs). Diffusion models similarly use a conditioned variational, and are similar to hierarchical VAEs (prince_understanding_2023, ). However, diffusion models do not use variational parameters $\phi$, and their latent variables have the same dimensionality as the observed variable, hence their use of $\bm{x}$ instead of $\bm{z}$.
@@ 3.4.2. VAEs and Diffusion Models
VAEs share many properties with PCNs, as discussed in (marino_predictive_2019, ; marino_predictive_2022, ; ofner_generalized_2022, ; pinchetti_predictive_2022, ). They make use of a similar graphical model as two-layer PCNs (and linear factor models, fig. 13). VAEs are trained using variational free energy as an upper bound for the NLL, cf. appendix A for a treatment of PCNs in these terms. While they use different optimization techniques, these can be seen as design choices for solving the same inference problem. A number of key differences can be identified. First, the conditional $p(\bm{x}|\bm{z})$ is parametrized by a linear function, followed by a nonlinearity (i.e.$f(W^{z}\bm{z})$). In contrast, VAEs use a deep neural network (trained with BP). Moreover, in the variational inference perspective, whereas PCNs use a variational posterior of the form $q(\bm{z})$, VAEs use $q_{\phi}(\bm{z}|\bm{x})$, i.e. with an explicit conditioning on data (visualized by dotted lines in fig. 13). This is parameterized by a separate neural network through parameters $\phi$, which are learned jointly with $\theta$ using BP. For further comparisons, see (marino_predictive_2019, ; marino_predictive_2022, ).
Adding multiple layers to VAEs results in hierarchical VAEs, which can similarly be compared to multi-layer PCNs. Diffusion models are also shown in fig. 13, which share (most of) their graphical structure with hierarchical VAEs. Like VAEs, diffusion models factorize the variational posterior using conditional distributions, i.e. $q(\bm{x}^{\ell}|\bm{x}^{\ell+1})$, but this is not parametrized by separate parameters $\phi$. I.e. the variational model is not learned, but fixed. These differences are summarized in table 7.
Table 7. Differences in the inference model $q$ between VAEs, diffusion models, and PCNs. Data dependence means $q(\bm{z}|\bm{x})$ instead of $q(\bm{z})$.
Fixed Explicitly data-dependent
(Hierarchical) VAE No Yes
Diffusion model Yes Yes
PCN No No
In terms of performance, comparisons between multi-layer PCNs and (non-hierarchical) VAEs have been studied by (ororbia_neural_2022, ; ofner_generalized_2022, ; pinchetti_predictive_2022, ; zahid_sample_2024, ). PCNs seem to compare favourably, with similar or improved results on several metrics, especially in the most recent work of (zahid_sample_2024, ). At the same time (similar to the picture that emerged for PCNs used for supervised learning), PCNs seem to be more computationally costly to train, though again relatively little work has been done on training optimization in comparison to that for VAEs. Comparisons with hierarchical VAEs and diffusion models have not yet been done to the authors’ knowledge.
@@ 3.4.3. Energy-Based Models (EBMs)
PCNs are often cited to belong to the class of EBMs (millidge_theoretical_2022, ; millidge_backpropagation_2022, ; song_inferring_2024, ). However, in these works the term ‘EBM’ is used in a different way than is typical in ML. The typical use of EBM (lecun_tutorial_2006, ; tomczak_deep_2022, ) is based on the observation that any probability distribution function can be parametrized by:
@@
where $E_{\theta}(\bm{x})$ is the energy function, and $Z(\theta)$ is the partition function, i.e., the sum of all possible states. This use comes from physics, where it is called a Boltzmann distribution, corresponding to a system at thermal equilibrium described by a canonical ensemble. The prototypical example of EBMs are Boltzmann machines, which have $E_{\theta}(\bm{x})=-\frac{1}{2}\bm{x}^{T}W\bm{x}$ (ackley_learning_1985, ; mackay_information_2003, ).
In contrast, in (millidge_theoretical_2022, ; millidge_backpropagation_2022, ; song_inferring_2024, ) an EBM is instead defined through (21) and (22) (millidge_theoretical_2022, ; millidge_backpropagation_2022, ; song_inferring_2024, ). This definition has also been used by (scellier_equilibrium_2017, ), and encompasses continuous Hopfield networks and networks trainable by Equilibrium Propagation. Indeed, PCNs have much in common with these techniques, cf. (millidge_backpropagation_2022, ) and section 4.
Regarding EBMs as defined by (34), comparisons of PCNs with this class of models has not yet been done to the authors’ knowledge. We briefly mention some differences here, comparing to Boltzmann machines. If one includes latent variables $\bm{z}$, one may write $p_{\theta}(\bm{x})=\sum_{\bm{z}}p_{\theta}(\bm{x},\bm{z})$ (with a sum instead of an integral because nodes in Boltzmann machines are typically discrete), with (mackay_information_2003, ):
@@
Notably, Boltzman machines are characterized by all-to-all connectivity between nodes. [^11] Another important class of models is obtained by restricting the weight matrix to only include connections between $\bm{x}$ and $\bm{z}$, whereupon one obtains restricted Boltzmann machines (RBMs), which can be stacked to obtain a hierarchical structure with interesting Bayesian interpretations (goodfellow_deep_2016, ). Now, observe that the generative model of PCNs, by definition of $E$ (cf. (20)), can be written as
(36) $p_{\theta}(\bm{x},\bm{z})=\exp(-E_{\theta}(\bm{x},\bm{z}))~{}.$
which is similar to (34). However, note that $E$ in this equation is not an ‘energy’ due to the lack of normalization in $Z$. Nonetheless, comparing EBMs to PCNs one may first observe that $E$ is different. Moreover, they are trained in different wat PCNs minimize $-\ln p_{\theta}(\bm{x},\bm{z})$, whereas EBMs typically minimize the NLL itself – not by using EM, but taking derivatives of the NLL w.r.t. model parameters, calculated using MCMC techniques, e.g. Gibbs sampling (goodfellow_deep_2016, ).
@@ 4. PCNs and Inference Learning
Where section 2 gave a general introduction to PCNs and IL, this section discusses IL in detail, which has been the focus of much of the recent literature. Most work in this area has focused the connection between IL and BP, introducing various ways of modifying IL to approximate or replicate BP’s weight updates. More recent papers have instead started considering IL in its ‘natural regime’ as presented in section 2. It has also become understood that PCNs/IL share properties with other algorithms, like target propagation (millidge_theoretical_2022, ) and trust region methods (innocenti_understanding_2023, ). We review these below. In general, much is still unclear about the properties of PCNs and IL. In particular, the properties of IL in generative PCNs and PC graphs remain underexplored.
@@ 4.1. Relation to BP
As far as we are aware, the first use of PC applied to an ML task was (whittington_approximation_2017, ). They showed, under certain strict conditions, that the weight updates of IL on discriminative PCNs approximates the parameter updates of BP in FNNs. Subsequently, this result was shown to generalize – again with strict conditions – to any computational graph (a decomposition of complex functions, such as complicated DL architectures, into elementary functions) (millidge_predictive_2022, ; rosenbaum_relationship_2022, ). Using a different method, it was shown that a variation of PC, called Z-IL, gives the exact weight updates of BP on any computational graph (i.e. including FNNs, CNNs and RNNs) (song_can_2020, ; salvatori_predictive_2021, ; salvatori_reverse_2022, ). The reason this works is that Z-IL uses a highly specific number of inference steps that matches the number of layers, and updates weights of different layers at very specific steps during inference. A requirement is that the ‘feedforward pass’ initialization of hidden nodes is done (see section 2.2.3). Alg. 4 shows pseudocode for Z-IL  [^12] Note that to write down Z-IL most simply we here use notation that contradicts the earlier sections. Here we use $\bm{\mu}^{\ell}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$, with data $\bm{x}^{(n)}$ clamped to $\bm{a}^{L}$ to and $\bm{y}^{(n)}$ to $\bm{a}^{0}$..
1:: $\bm{\mu}=f(\bm{w}^{\ell+1}\bm{a}^{\ell+1})$ // Upward predictions
2:: $\bm{a}^{L}=\bm{x}^{(n)}$, $\bm{a}^{0}=\bm{y}^{(n)}$ // Clamp data
4:: $\gamma=1$
5:for $t=0$ to $L$ do // Note: $T=L$
6:   for each $\ell$ do
9:   end for
10:end for
@@
Since it has been known that other algorithms (Equilibrium Propagation (scellier_equilibrium_2017, ), Contrastive Hebbian Learning (xie_equivalence_2003, )) also approximate BP, work by (millidge_backpropagation_2022, ) showed that these various approximations can be understood as a consequence of certain properties of energy-based models (EBMs), as defined by (21) and (22) (cf. section 3.4.3). A practical limitation of these works is that the various approximations and variants of PC always remain slower than the (highly optimized) BP (zahid_predictive_2022, ). As such, they do not provide a clear argument for using e.g. Z-IL in practical applications.
@@
Another interesting theoretical connection found by (millidge_theoretical_2022, ) is that PC can be understood to interpolate between BP and a third algorithm for training neural networks, called target propagation (TP) (lee_difference_2015, ). This is algorithm propagates target activations (not errors) to the hidden layers of the network, followed by updating weights of each layer to move closer to the target activation. The authors show that if a PCN is trained with only labels clamped, the activies converge to the targets computed by TP. At the same time (as was shown in section 2), with only data clamped, the activations computed with IL are equal to feedforward ANN activations. As such, the equilibrium of IL’s inference phase with both data and labels clamped can be interpreted as an average of the feedforward pass values of the network, and the local targets computed by TP. Furthermore, they observe that precision weights $\Pi^{\ell}=(\Sigma^{\ell})^{-1}$ provide a means to change the relative weighting of the feedforward and feedback influences on the network dynamics (cf. section 3.3.2).

### v2

@@ 3.2.2. Multi-layer PC
As in the simple model above, we assume a multidimensional Gaussian for the conditional distribution of each layer, as well as for the prior: [^13] This is a reasonable assumption for traditional (fully-connected, feedforward) ANNs with large layer widths, $n_{\ell}\sim n\gg 1$, which becomes exact in the so-called large-width limit $n\mapsto\infty$, due to the central limit theorem. However, non-Gaussianities appear at finite width (see, e.g., (Roberts et al., 2022; Grosvenor and Jefferson, 2022) for theoretical treatments), and it is unclear how realistic this assumption is in more general networks involving intralayer or feedback connections.
@@
which is the energy of Section 2; cf. also (26). Applying EM to this then yields IL as presented in section 2 (replacing $\bm{z}^{\ell}$ by $\bm{a}^{\ell}$ [^14] Following typical ML notation, we use $\bm{z}$ for latent variables in the context of graphical models, and $\bm{a}$ for activation neurons in the neural network context.).
@@ 3.3.1. Discriminative and Generative PCNs
At this point we can provide a number of interesting interpretations of earlier sections. Notice that in the derivation above, the model had observed variables at layer $\ell=0$. This corresponds to generative PCN, as discussed in section 2.5. If trained in supervised mode, we can understand the clamping of the label $\bm{y}^{(n)}$ to the final layer $\bm{z}^{p}$ as follows: $\bm{y}^{(n)}$ becomes a parameter of the prior (e.g. its mean). If trained in unsupervised mode, it is treated just like the other hidden variables. Interestingly then, for a discriminative PCN, we see that the clamped data $\bm{x}^{(n)}$ is a parameter for the prior, and the label is now the observed variable. This is shown in table 6.
Table 6. Overview of learning modes in the probabilistic perspective of PCNs.
@@ 3.3.2. Precision Matrices
In the context of predictive coding, the covariance matrices $\Sigma^{\ell}$ are often interpreted as the uncertainty at each layer around the mean $\bm{\mu}^{\ell}$, with their inverse $(\Sigma^{\ell})^{-1}\eqqcolon\Pi^{\ell}$ sometimes called precision matrices (Friston, 2003; Millidge et al., 2023b; Marino, 2022). Within neuroscience and psychology, they have been used in a wide range of models (Millidge et al., 2022a), where it has been suggested that they implement a type of attention (Feldman and Friston, 2010). By up-weighting the error, increasing the precision of a variable would serve as a form of gain modulation. From the ML perspective, instead of setting them to identity matrices, one can add minimization of $E$ w.r.t. $\Pi^{\ell}$ as an additional part of the M-step, seeing them simply as additional parameters of the generative model. However, their practical utility remains unclear, with varying results. In particular, (Ofner et al., 2022) do not find advantages of additionally learning for generative modeling tasks, while (Ororbia and Kifer, 2022) do appear to find some benefit, suggesting they act as a type of lateral modulation that induces a useful form of sparsity in the model. Comparisons with VAEs are made in (Pinchetti et al., 2022), with a different theoretical interpretation by (Millidge et al., 2023b) (cf. section 4 for further discussion).
@@ 3.4. Connections to Other Latent Variable Models
With $f$ as a linear function, the two-layer PCN model discussed above is equivalent to a linear factor model. This is illustrated in fig. 13. Specifically, if we take $\bm{\mu}_{z}=\bm{0}$, $\Sigma_{z}=I$, and $\Sigma_{x}=\operatorname{diag}\bm{\sigma}^{2}$ with $\bm{\sigma}^{2}=[\sigma_{1}^{2},\sigma_{2}^{2},...,\sigma_{n}^{2}]^{T}$, then we obtain exactly factor analysis. If additionally take $\Sigma_{x}=\sigma^{2}I$, i.e. the variances $\sigma_{i}^{2}$ are all equal to each other, we obtain probabilistic PCA (Goodfellow et al., 2016). Like in PCNs, parameters in this model can be learned using the EM algorithm (Bishop, 2006).
Figure 13. Graphical model comparison of PCNs and other probabilistic models, with $\theta$ model parameters and $\phi$ variational parameters (cf. appendix A). Standard PCNs (left column) use MAP inference of the latent variables during the E-step, without variational parameters. Using a Gaussian variational (middle column) adds the mean and variance of the variational as parameters $\phi$ which may additionally be varied, shown by dotted lines. This is similar to $\phi$ in a VAE (top right). VAEs differ from PCNs in that their variational posterior is conditioned explicitly on $\bm{x}$ (i.e. $q(\bm{z}|\bm{x})$ in VAEs vs. $q(\bm{z})$ in PCNs). Diffusion models similarly use a conditioned variational, and are similar to hierarchical VAEs (Prince, 2023). However, diffusion models do not use variational parameters $\phi$, and their latent variables have the same dimensionality as the observed variable, hence their use of $\bm{x}$ instead of $\bm{z}$.
@@ 3.4.2. VAEs and Diffusion Models
VAEs share many properties with PCNs, as discussed in (Marino, 2019, 2022; Ofner et al., 2022; Pinchetti et al., 2022). They make use of a similar graphical model as two-layer PCNs (and linear factor models). VAEs are trained using variational free energy as an upper bound for the NLL (cf. appendix A). While they use different optimization techniques, these can be seen as design choices for solving the same inference problem (Marino, 2022). A number of key differences can be identified. First, the conditional $p(\bm{x}|\bm{z})$ is parametrized by a linear function, followed by a nonlinearity (i.e.$f(W^{z}\bm{z})$). In contrast, VAEs use a deep neural network (trained with BP). Moreover, in the variational inference perspective, whereas PCNs use a variational posterior of the form $q(\bm{z})$, VAEs use $q_{\phi}(\bm{z}|\bm{x})$, i.e. with an explicit conditioning on data (visualized by dotted lines in fig. 13). This is parameterized by a separate neural network through parameters $\phi$, which are learned jointly with $\theta$ using BP. For further comparisons, see (Marino, 2019, 2022).
Adding multiple layers to VAEs results in hierarchical VAEs, which can similarly be compared to multi-layer PCNs. Diffusion models are also shown in fig. 13, which share (most of) their graphical structure with hierarchical VAEs. Like VAEs, diffusion models factorize the variational posterior using conditional distributions, i.e. $q(\bm{x}^{\ell}|\bm{x}^{\ell+1})$, but this is not parametrized by separate parameters $\phi$. I.e. the variational model is not learned, but fixed.
In terms of performance, comparisons between multi-layer PCNs and (non-hierarchical) VAEs have been studied by (Ororbia and Kifer, 2022; Ofner et al., 2022; Pinchetti et al., 2022; Zahid et al., 2024). PCNs seem to compare favourably, with similar or improved results on several metrics, especially in the most recent work of (Zahid et al., 2024). At the same time (similar to the picture that emerged for PCNs used for supervised learning), PCNs seem to be more computationally costly to train, though again relatively little work has been done on training optimization in comparison to that for VAEs. Comparisons with hierarchical VAEs and diffusion models have not yet been done to the authors’ knowledge.
@@ 3.4.3. Energy-Based Models (EBMs)
PCNs are often cited to belong to the class of EBMs (Millidge et al., 2023b, a; Song et al., 2024). However, in these works the term ‘EBM’ is used in a different way than is typical in ML. The typical use of EBM (Lecun et al., 2006; Tomczak, 2022) is based on the observation that any probability distribution function can be parametrized by:
@@
where $E_{\theta}(\bm{x})$ is the energy function, and $Z(\theta)$ is the partition function, i.e., the sum of all possible states. This use comes from physics where (33) is called a Boltzmann distribution, corresponding to a system at thermal equilibrium described by a canonical ensemble. The prototypical example of EBMs are Boltzmann machines, which have $E_{\theta}(\bm{x})=-\frac{1}{2}\bm{x}^{T}W\bm{x}$ (Ackley et al., 1985; MacKay, 2003).
In contrast, in (Millidge et al., 2023b, a; Song et al., 2024) an EBM is instead defined through (21) and (22) (Millidge et al., 2023b, a; Song et al., 2024). This definition has also been used by (Scellier and Bengio, 2017), and encompasses continuous Hopfield networks and networks trainable by Equilibrium Propagation. Indeed, PCNs have much in common with these techniques, cf. (Millidge et al., 2023a) and section 4.
Regarding EBMs as defined by (33), comparisons of PCNs with this class of models has not yet been done to the authors’ knowledge. We briefly mention some differences here, comparing to Boltzmann machines. [^15] Boltzman machines are characterized by all-to-all connectivity between nodes. Another important class of models is obtained by restricting the weight matrix to only include connections between $\bm{x}$ and $\bm{z}$, whereupon one obtains restricted Boltzmann machines (RBMs), which can be stacked to obtain a hierarchical structure with interesting Bayesian interpretations (Goodfellow et al., 2016). If one includes latent variables $\bm{z}$, one may write $p_{\theta}(\bm{x})=\sum_{\bm{z}}p_{\theta}(\bm{x},\bm{z})$ (with a sum instead of an integral because nodes in Boltzmann machines are typically discrete), with (MacKay, 2003):
@@
Now, observe that the generative model of PCNs, by definition of $E$ (cf. (20)), can be written as $p_{\theta}(\bm{x},\bm{z})=\exp(-E_{\theta}(\bm{x},\bm{z}))$, which is similar to (34). However, note that $E$ in this equation is not a proper ‘energy’ as in (33) due to the lack of normalization in $Z$. Comparing EBMs to PCNs, one may first observe that $E$ is of course different. Another difference is that PCNs minimize $-\ln p_{\theta}(\bm{x},\bm{z})$, whereas EBMs typically minimize the NLL itself – not by using EM, but taking derivatives of the NLL w.r.t. model parameters, calculated using MCMC techniques, e.g. Gibbs sampling (Goodfellow et al., 2016). Clarifying these differences further could be an interesting direction for future work, further bringing PC and ML closer together.
@@ 4. PCNs and Inference Learning
This section discusses IL in detail, which has been the focus of much of the recent literature – particularly on the connection between IL and BP. Various ways of modifying IL have been introduced to approximate or replicate BP’s weight updates. More recent works however, instead consider IL in its ‘natural regime’ – meaning without modifications as presented in section 2. It has also become understood that PCNs/IL share properties with other algorithms, like target propagation (Millidge et al., 2023b) and trust region methods (Innocenti et al., 2023). We review these below. In general, much is still unclear about the properties of PCNs and IL. In particular, the properties of IL in generative PCNs and PC graphs remain underexplored.
@@ 4.1. Relation to BP
As far as we are aware, the first use of PC applied to an ML task was (Whittington and Bogacz, 2017). They showed, under certain strict conditions, that the weight updates of IL on discriminative PCNs approximates the parameter updates of BP in FNNs. Subsequently, this result was shown to generalize – again with strict conditions – to any computational graph (a decomposition of complex functions such as DL architectures, e.g. FNNs, CNNs and RNNs, into elementary functions ) (Millidge et al., 2022a; Rosenbaum, 2022). Using a different method, it was shown that a variation of PC, called Z-IL, gives the exact weight updates of BP on any computational graph (Song et al., 2020; Salvatori et al., 2021b, 2022b). The reason this works is that Z-IL uses a highly specific number of inference steps that matches the number of layers, and updates weights of different layers at very specific steps during inference. Another requirement is that the ‘feedforward pass’ initialization of hidden nodes is done (cf. section 2.2.2). Alg. 4 (appendix B.4) shows pseudocode for Z-IL.
@@
Since it has been known that other algorithms (Equilibrium Propagation (Scellier and Bengio, 2017), Contrastive Hebbian Learning (Xie and Seung, 2003)) also approximate BP, work by (Millidge et al., 2023a) showed that these various approximations can be understood as a consequence of certain properties of energy-based models (EBMs), as defined by (21) and (22) (cf. section 3.4.3). A practical limitation of these works is that the various approximations and variants of PC always remain slower than the (highly optimized) BP (Zahid et al., 2023a). As such, they do not provide a clear argument for using e.g. Z-IL in practical applications.
@@
Another interesting theoretical connection found by (Millidge et al., 2023b) is that PC can be understood to interpolate between BP and a third algorithm for training neural networks, called target propagation (TP) (Lee et al., 2015). This is algorithm propagates target activations (not errors) to the hidden layers of the network, followed by updating weights of each layer to move closer to the target activation. The authors show that if a PCN is trained with only labels clamped, the activies converge to the targets computed by TP. At the same time (as was shown in section 2), with only data clamped, the activations computed with IL are equal to feedforward ANN activations. As such, the equilibrium of IL’s inference phase with both data and labels clamped can be interpreted as an average of the feedforward pass values of the network, and the local targets computed by TP. Furthermore, they observe that precision weights $\Pi^{\ell}=(\Sigma^{\ell})^{-1}$ provide a means to change the relative weighting of the feedforward and feedback influences on the network dynamics (cf. section 3.3.2).